Information service via the Internet: the present and the future - Metaphysical issues and JPNIC -

Haruyuki Tatsumi

Department of Anatomy, Sapporo Medical University School of Medicine, Sapporo, Japan E-mail:tatsumi@sapmed.ac.jp A member of the database committee of Japanese Association of Anatomists Manager of Japan Committee for Research Networks (JCRN)[1-3] A member of the steering committee of Japan Network Information Center(JPNIC)[4] Representative of Network Organization of Research and Technology in Hokkaido (NORTH)[5]

1 Introduction: A HyperEnd User?

I am an anatomist using computer workstations for various purposes. Is it so attractive? I do not know why, but, I am often solicited to write some articles or to make some lectures on computers[6-12]. To my surprise, I was invited to a computer specialist meeting in order to talk about computer utilization. As is often the case with the specialists, they asked me out of curiosity, “For what are you using computers?” They expected me to use them for a special use in anatomy. To tell the truth, that is my first intention[13-24], however, I am making use of workstations for various things as I understand what computers are. We, in the medical field, have tons of information up to our neck and need more special medical information than that. Since we are under the necessity of dealing with a lot of information, we have to more or less take advantage of computers and networks. As a hyperend user, I would like to make some comments on the background of computer and its utilization and refer to the Internet information service, including the Japan Network Information Center (JPNIC)[4] activities.

2 Backgrounds in Computer Utilization

2.1 A comparison between mankind and computer in history

• Mankind created languages.

One could begin to communicate with the other persons.

• Mankind produced characters and paper. Ideas and events could be recorded and passed on to the next generation.

• Mankind invented movable type printing.

This enabled us to write more and more legible and eventually led to the creativity of books, one of man’s greatest feat. The beginning of the spread of civilization.

• Mankind developed message delivering system.

For example, postal service, telecommunication and computer networks.

Similar achievements are found in the developmental history of computer.

• Mankind created computer languages and operating systems (OS). This enabled us to control computers by using words (not with 1 and 0).

• Mankind invented many kinds of recording media for digital data.

Magnetic tapes and floppy, hard and optical disks. These are being substituted for paper.

• Mankind made word processors and desk top publishing software.

Our idea or concept can be processed and printed by using computers.

• Mankind developed computer networks.

Following the same developmental processes, computers are changing their pictures from a simple calculator to various tools for intellectual activities, coming close to us and accelerating our development.

2.2 Many aspects of computers

Pictures about computers in our minds are totally different from each other, because computers can serve as various machines according to software. This sometimes throws us into disagreement when we are talking about computers. I will name some of the aspects.

• Family computers (so called ”fami-com” in Japan).

• TV games on personal computers.

• Word processors (”wor-pro” in Japan, which is a special computer for editing and printing).

• Computer networks.

• Data base systems.

• Multivariate statistical analysis, scientific calculation and solution of arithmetic problem.

• Expert system, artificial intelligence and neural nets.

• Data processing (signals, numerals, characters, images and so on)

There exists so many usages as computer programs are developed. The more useful they are, the more complicated the way to use is. That is why software developers are making more user-friendly software so that users find them less complicated.

3 Issues in Computer Utilization

What is the best computer? The crucial point is the speed, the number of functions or the simple user-interface? It is up to the users. Anyway, simple is the best, of course, but there is a pitfall.

3.1 Abuse of dedicated machines in Japan

Workers are busy with their jobs, so they would like to use computers to save their time. They have to make time to learn how to use, so they have little room to do extra things, i.e. studying computer hardware, OS and programming. Japanese companies noticed the business chances, giving them much convenience with the specialized computers. Therefore, ”wor-pro” and ”off-com (office computers)” are thriving in Japan. This is a good thing in the short run, but a mischance in the long run. The users lost the chance to study computers and software under pressure of necessity.

Analogy-people use languages today in every culture. It is not really necessary for them to know how the language was developed for them to use it effectively as tools and in communications in their every day lives. If the intellectual or other interested person wants to study the characters or developmental history, then they are free to do this.

3.2 Difficult and easy to use

Computers are originally calculators, however, it is very hard to calculate with personal computers, workstations or main frames. Do you think so? These computers need special programs for calculation. As for the workstation, we must start up the machine with the OS (UNIX), login the system, and run the X Window System, a window manager and a calculator program. There are complicated steps, which deter us from using computers. But once we set the procedures in a program, it is very easy to use. Therefore, we had better learn sort of programming to some extent. If someone set up some procedures in a program in stead of you, you can use it with ease, but not at will. At that case, it is very difficult to modify the program unless you know how to set up.

Computers have great potential, changing into assistant tools for intellectual activity. Their value much depends on the software and much more on the software’s software that is in our mind. Therefore, we have to cultivate the ability to use software. In a sense, the dedicated computers spoil us. Dedicated computers have specific uses, therefore average people buying computers should be aware of this and buy computers that offer more versatile use with more useful software.

3.3 Different feeling

In terms of what we can do with computers, the critical point is quite different between engineering and medical fields. I asked some of the professors of engineering about some programs, “Can you make?” They would often say, “Yes, we can of course.” It was a very kind of them. I immediately begged them to use it. They were usually at a loss to make an excuse. They said, “It is certainly possible if we try to make it. But, we haven’t tried yet. It takes a long time. I am sorry.” I have noticed that their sense of time is very different. If it is possible, I want to do it now. We, in medical field, cannot wait so long. Many things are theoretically possible with computers, but it is usually difficult to do immediately.

Dedicated software engineers can and often will create the necessary software to do and solve most problems within the scope of their scientific knowledge if the task is properly specified; therefore if medical doctors are able to describe their need, there are competent software engineers who can write the necessary software - some times to their own surprise, they shock themselves. To describe our need properly, we had better know about computers and programs.

Productivity of the software should be improved. If they can make it in a few minutes, they would do it for me. Because it takes substantial days or man-hours, they would not accept my entreaty. In addition, much more important things are motive power which let them overcome the crucial point between “to do” and “not to do”. There is an episode: My friend had decided to buy a computer. But it took two years to buy it without any reason. I can not think he had decided. He had been hesitating for two years!

3.4 Contradictions in computerization and the Internet

In order to meet the users’ demands, computer programs consist of so many functions that it is impossible to know all of them. We can not apply the computer to our job unless we know a little bit about the computer and the program in advance. It is very difficult to learn them practically unless we possess either of them. Somehow, we have to take a chance to introduce computers.

This is the case with the Internet, which was popular only in the limited number of disciplines. Nowadays the Internet is indispensable and so widely and rapidly spreading out to be an environment. If you do not have any connectivity to the Internet, it is hard to understand the usefulness of the environment.

3.5 Computer literacy

Literacy means abilities to read and write. In Japan, from old days, similar standard abilities learned in a “terakoya (temple school in old days)” have been “yomi(read) kaki(write) and soroban(calculation).” Besides so-called literacy, calculation had been a standard requirement in Japan. I am proud of the fact, inasmuch as Japan is far behind the United States in computer utilization but was a step ahead in old days.

At first, computers played the part of calculation. Then, due to the advent of word processors and so on, computers came to be used for the literacy part. So the ability to utilize computers is the literacy for the next generation. The computers and networks are of extensive use. We had better cultivate the ability to use them as soon as possible.

3.6 Other issues in computer utilization I name some of the issues, which I do no refer to in detail here.

• Paper less or more with computer utilization?

• Paperwork is very important. No job is finished till the paperwork is done!

• Distinct purpose is needed.

• Front end processor for Japanese.

• Shortage in computer technologists having a special line other than computer fields.

• English proficiency is very important: Scanty of manuals and documentation in Japanese.

• The importance of the trifle information for computer utilizations.

• What is the best computer: many functions or a simple keyboard?

• Computer dependent syndrome is important?

• Fail-safe and backup system, just in case.

4 Some of the solutions for the issues

Isolated computers and fragmented networks are being integrated into a worldwide network, the Internet. Thus, human progress is much more accelerated by computer utilization. Therefore, we had better to take advantage of them, even to solve the computer problems themselves. First we have to use computer for our purposes, i.e. personal computing.

4.1 Personal computing For what can researchers use computers? Except for the specialists in computer science, most of the researchers are not professional in computers. As I described above, computers are undergoing metamorphosis into intellectual tools and information processing machine. Therefore, computers come to have lots of possibilities in many fields. Then, how can personal computing assist our activity? On the contrary, what is it that we need in common?

• Integrated information system.

– Acquisition of information

– Management (Reuse) of information

– Exchange of information

• Data processing and analysis system.

• Publishing system.

– Tools for writing dictionary, thesaurus, lexicon, spell checker, grammar checker, idea processor, outline processor, reference maker

– Tools for visualization (image, graph, schema, figure, etc.)

– Tools for printing (DTP:Desk Top Publishing)

For comfortable personal computing, various tools must be integrated with each other and connected to the Internet. This is important not only for the personal computing but also for the interpersonal communication. Nowadays, some of the scientific data should be submitted as computer readable data, which save time and money to input and correct the data. If you use computers connected with the Internet in order to make the data, it is very easy to send the data via e-mail.

4.2 Interpersonal communication

Though research activities are principally based on personal activity, the interpersonal communication have a great influence on research activities. In addition, the result of the research should be published and reused by others. The personal computing assists the personal activity as well as facilitates the interpersonal communication. Reversely, the enhanced interpersonal communication will solve some of the problems in personal computing. To get such a comfortable environment, the following points are recommended as countermeasures.

• Use the same software.

• Unify the user interface of software.

• Develop the network community where we study and teach each other.

One of the answers is GNU software, which is distributed over the Internet. In the GNU manifest, the general public license is intended to give everyone the right to share GNU’s software. To make sure that we get the rights they want us to have, they need to make restrictions that forbid anyone to deny us these rights. GNU software facilitate both the personal computing and the interpersonal communication, leading to the development of the Internet.

5 Backgrounds in the Internet

In general, we are not usually aware of “an environment.” For example, we do not take care of the air. If the air pollution had happened or the air had disappeared, we would puzzle and notice that the air was very important. We do not find the environment is important until we lose it. Since the environment itself is not so useful, it is difficult to do it justice. But it becomes indispensable once we come to utilize it. The Internet in Japan has not been an environment for everyone yet. We are just at the very beginning of the Internet as an environment.

5.1 An environment : the Internet

Good research environments are not optional, but a standard of living in colleges and universities. As an infrastructure of academies, computer and network systems should be installed like electric supply or water supply. Local area network (LAN) is not useful in fact if there is no connection with outside. The network should be interconnected with local, regional and wide area networks. In Japan, campus LANs are being installed in many universities and colleges. In this way, the Internet in Japan started from research fields.

Networks must be installed on a large scale, like the Internet, which is almost on an environment level. Localized computerization and fragmented networks are not so powerful. Under threshold, it is almost nothing by the all-or-none law. Halfway is bad. Since the NII (National Information Infrastructure) project in the United States is big, it is worth while executing. Though Japan is far behind the United States in the computerization, Japanese government is going to execute a similar project in a shorter period as compared with the United States. Before that, we need a revolution in our mind and have to be able to deal with the Internet.

5.2 Issues in the Internet

There are very few consultant companies which can help you to introduce the Internet into your office, in Japan. Needless to say, the consultant company must master the Internet.

Let me give an example at hand. In my university, an anatomist (it’s me) has to administrate and maintain the local area network and the connection to the Internet. That’s a very strange thing, but I can not help t and have to use the Internet. We do not have any technologist and such kind of position in our school. This kind of thing could be applied to many other companies and departments in Japan. A typical example is the library, which is forced to use computer and the Internet. Physical things are easy to see and understand. Seeing is believing. It is apparently not so difficult to install computers and the Internet. But the metaphysical part is difficult to understand and has to be revolutionized.

The Internet is so novel that anyone can not forecast the future. It has extra-ordinal potential and problems at the same time. In Japan, some of the administrative work for the Internet was carried out by volunteers, making an effort to eliminate the volunteer work to facilitate the development of the Internet.

5.3 The genesis of JCRN and JPNIC

With the rapid growth of the Internet in Japan, many people needs correct information about the Internet. Moreover, most important things are official assignment and registration for domain names and network numbers, which are guaranteed to be unique all over the world. Without the administrative work, which was carried out by voluntary group, the Internet could not have survived in Japan.

As I mentioned above, the Internet is so useful especially in academic fields that the retardation of the development of the Internet is a stumbling block to the progress of the science. In order to coordinate various issues concerned with research networks, the Japan Committee for Research Networks (JCRN) was established on October 16, 1990. The aim is stated in a document (anonymous ftp:nic.ad.jp/pub/jcrn/RULE.txt). JCRN consists of the delegates from scientific societies and associations and research networks, which have a concern in computer communication. In JCRN, the necessity of a network information center for Japanese computer networks was discussed.

For outgrowth from the conventional voluntary system of volunteers of each network organization to a system for fair and rapid processing by the cooperation of each network project, the Japan Network Information Center (JNIC) was inaugurated in December 1991 under JCRN, and JNIC has kept activity for about one and half years. In consequence, allocation and administration of network addresses, a system for providing the network information and estimation of work quantity have become possible. Besides, JNIC has been used as a service though not satisfactory. However, organizational foundation, especially financial foundation, could hardly be improved. Therefore, it is hard to establish a system sufficient for supporting the development of computer networks and to correspond to commercial internets. Under these circumstances, the Japan Network Information Center was newly established as an independent organization from JCRN. The new Japan Network Information Center is abbreviated as JPNIC, for relations with other countries.

6 JPNIC

• Roles:

– IP network number assignment

– JP domain name allocation

– JP name server administration

– JPNIC database maintenance

– Information services (see below)

– Cooperation with other domestic organizations (JCRN, JEPG, etc.)

– International cooperation (InterNIC, RIPE NCC, APNIC[26], etc.)

• Rules: As defined by the rules of the Japan Network Information Center (see anonymous ftp:nic.ad.jp/pub/jpnic/jpnic-rule.txt), this Center was established as a voluntary organization. The committee is planning to incorporate JPNIC.

• Secretariat:

c/o Computer Centre, University of Tokyo 2-11-16, Yayoi, Bunkyo-Ku, Tokyo 113 Japan

• Establishment Date: March 10, 1993 6.1 The JPNIC information services

Cited from anonymous ftp:nic.ad.jp/pub/jpnic/jpnic-services.txt

      Brief Guide to JPNIC Information Services
     (May, 1993)
Japan Network Information Center (JPNIC)
c/o Computer Centre, Univ. of Tokyo
Yayoi 2-11-16, Bunkyo-ku, Tokyo 113, Japan

JPNIC is a country NIC of Japan. All information of JP domain names,
IP network numbers, name servers, contact persons, and network
providers in Japan are collected into JPNIC database to be retrieved
via whois service. JPNIC cooperates with InterNIC to keep the database
consistent.
Here is an example to retrieve the JPNIC database with whois. If you
want to know what domain name "ABC University" has, execute a whois
command on your machine as follows:
% whois -h whois.nic.ad.jp ABC/e
It may produce a lot of summary lines including other information such
as IP addresses but you will find its domain name if it has. After
getting the domain name ABC.AC.JP, you'd better check its full content
as follows (if "ABC" matches only single entry, whois generates its
full content directly):
% whois -h whois.nic.ad.jp ABC.AC.JP/e
If its state is not "Connected" you do not reach at the domain. Even if
the state is "Connected" some domains are not reachable from outside of
Japan depending on its right of international link use.
JPNIC whois has similar functions to InterNIC whois. If you want know
more on its usage please send help as follows:
% whois -h whois.nic.ad.jp HELP/e
If your machine doesn't have whois, use telnet instead as follows:
% telnet whois.nic.ad.jp 43
[ connecting messages... ]
ABC/e
If you unfortunately don't have direct Internet connectivity, you can
send a request via electronic mail as follows:
% mail mail-server@nic.ad.jp
whois ABC/e
If you can or want to read them in Japanese, omit trailing "/e" in any
case.
JPNIC database is for network operation. It includes "point of contact"
of the information but is NOT intended to be a general directory.

(Don't ask us your friend's address, please.)
All of the domain names under JP are also obtained from JPNIC via two
methods, anonymous ftp and electronic mail as follows:
% ftp ftp.nic.ad.jp
Name: ftp
Password: your_login_name
ftp> get pub/jpnic/domain-list-e.txt
% mail mail-server@nic.ad.jp
send jpnic/domain-list-e.txt
In the list domains not connected are enclosed by parentheses. If you
can't find a name in the list (as well as whois), it doesn't have a
domain name under JP. Most universities of BITNET Japan don't use JP
domain names yet so they are reachable via a gateway but don't appear
in the list. JPNIC database only maintains the third level of JP
domain names, so you can't find any organizations hidden under the
third level of a JP domain.
          COPYRIGHT NOTICE OF JPNIC OPEN DOCUMENTS
This is a JPNIC open document and all copyrights of this
document are with the Japan Network Information Center.
Anybody can obtain a JPNIC open document from JPNIC by
merely paying the mailing costs.  Further, anybody can
freely reproduce, copy, or redistribute this document
provided this copyright notice is also reproduced at the
same time.

7 The Internet information services

Thanks to basic information services, such as the JPNIC services, we can use the following services of the Internet: e-mail (mailing list), net-news, anonymous FTP, archie, gopher, www (xmosaic), irc (internet relay chat) and so on. I describe two of them.

7.1 World Wide Web (WWW) - xmosaic (Sapporo LAMeN server)

The World Wide Web is a wide area client-server architecture for retrieving hypermedia documents across the Internet. It also supports a means for searching remote information sources, for example bibliographies, phone directories and instruction manuals. The protocols include FTP, NNTP, WAIS, Gopher and HTTP. The latter is designed specifically for use with the World Wide Web, and combines efficiency with an ability to flexibly exchange information between clients and servers. One of the client programs is a xmosaic which was developed by NCSA (National Center for Supercomputing Applications). Figure 1 shows xmosaic windows, showing the Sapporo LAMeN server menu. From the menu, you can see the data of the State Examination for Medicine, go into the gopher server of the SAINS (SApporo medical university Information Network Systems) and so on. We can retrieve the data of the

fig 1. xmosaic, showing the sapporo LAMeN server of sapporo medical university. Fig 2. The zircon, a IRC client, showing a list window, two chat windows and a conversation window.

proceedings of Japanese Association of Anatomists, which were experimentally collected via the Internet[3, 9, 25].

7.2 IRC (Internet Relay Chat)

Figure 2 shows a client program of internet relay chat, Zircon. In the IRC channel list, there are lots of the channels which have each topic. The Zircon information window shows the name of IRC server and the number of persons who are using IRC. This is a kind of “talk” program in UNIX. This program enable us to talk with many persons on the Internet at the same time. Therefore it is very easy to get answers from many persons. In addition, we can discuss privately. Figure 2 shows two channels (#ちゃっと, #kyutech), and one private channels (IRC Conversation).

8 The Future in the Internet

Scientific research networks are the origin of the Internet in Japan. As mentioned above, JCRN played a valuable role in the development of the Internet, giving rise to the generation of JPNIC. Thereafter, JCRN is expected to facilitate research activities through computer communication in conjunction with the Ministry of Education, Science and Culture, other organizations and so on.

For the wealthy and healthy development of the Internet in the future, which is worth for not only research but also business fields, JPNIC should be reinfored from all quarters. This is a basic part of the Internet, on which the other information services of the Internet are dependent. Therefore we need first such a stable infrastructure and then we can pile up our efforts and fruits for the future.

Though physical things are important, there are many chances to improve them as compared with the metaphysical things, which are difficult to discern. From now on then, more and more of computers and networks will be installed in many places. We, users, have to be much more enlightened to adapt ourselves to the new environment and have to develop new common sense and morals in the computer communications.

With infrastructures alone, we can not attain great achievements. Metaphysical things, such as software’s software, the integration of the system and so on, are important to build up on the infrastructure. Paying attention to the metaphysical part of the Internet, I would like to play a medical role in the healthy develpment of the Internet as a hyperend user.

Acknowledgements

SAINS in Sapporo Medical University, which I introduced here, is maintained and improved by a voluntary group (LAN wo kangaerukai) in cooperation with faculties, students and office workers. I thank all of them for their support and Ms. Mai Luong at the University of Alberta, Canada, and Mr. Howie Rowe who are so kind to check this paper by using e-mail, talk and IRC.

References

[1] JCRN newsletter (in Japanese). anonymous ftp:nic.ad.jp/pub/jcrn newsletter-1.1.txt Vol. 1, No. 1, January (1991) newsletter-1.2.txt Vol. 1, No. 2, March (1991) newsletter-1.3.txt Vol. 1, No. 3, September (1991) newsletter-2.1.txt Vol. 2, No. 1, Decembrer 4, (1992).

[2] Tatsumi H. Research activity and computer network:Japanese Association of Anatomists participated in Japan Committee for Research Networks(JCRN). Acta Anat Nippon, 68:231-234 (1993) (in Japanese).

[3] Tatsumi H. Activities of the academic societies and JCRN (Japan Committee for Research Networks): the activities using the Internet. J Clin Exp Med (IGAKU NO AYUMI), 169:in press (1994) (in Japanese).

[4] JPNIC newsletter 1:1-49 (1994) (in Japanese).

[5] Kitada Y, Tatsumi H. The Network Organization of Research and Technology in Hokkaido (NORTH). Proceedings of WINC Network Symposium, 1-6 (1993) (in Japanese).

[6] Tatsumi H. Direction and prospect of computer utilization in biomedical science :Aim at comfortable research environment. J Clin Exp Med (IGAKU NO AYUMI), 152:107(1990) (in Japanese).

[7] Tatsumi H. ”Comfortable research environment”:academy equipment from personal computing to interpersonal communication via computer network. Proceedings of the 6th Sapporo International Computer Graphics Symposium, 12-14 (1992).

[8] Tatsumi H. Analogy between computers and living organisms: from a view point of anatomist. Electronic Network Consortium,4:12-13 (1993) (in Japanese).

[9] Tatsumi H. Application of the Internet to the Japanese Association of Anatomists. HINES World, 24:1-10 (1994) (in Japanese).

[10] Tatsumi H. Common property of the intellectual product and the Internet. めでぃあ (Media), 31:14(1994)(in Japanese)

[11] Tatsumi H. The present circumstances of the computer utilization and the Internet. J Clin Exp Med (IGAKU NO AYUMI), 169:1131-1134 (1994) (in Japanese).

[12] Tatsumi H. Problems in computer utilization. J Clin Exp Med (IGAKU NO AYUMI), 170:in press (1994) (in Japanese).

[13] Tatsumi H, Takaoki E, Kurosawa T, Omura K, Fujita H. A new reconstruction method from serial sections by computer graphics using ”meta-balls.” Proceedings of the 2nd Sapporo International Computer Graphics Symposium, 69-72 (1988).

[14] Tatsumi H, Takaoki E, Omura K, Fujita H. A new method for three-dimensional reconstruction from serial sections by computer graphics using ”meta-balls” : Reconstruction of ”Hepatoskeletal system” formed by Ito cells in the cod liver. Comp Biomed Res, 23:37-45(1990).

[15] Kakinuma R, Takaoki E, Tatsumi H. Three dimensional visualization of the trachea and main bronchi in tracheobronchomegaly by computer graphics using ”meta-balls.” Proceedings of the 3rd Sapporo International Computer Graphics Symposium, 31-35 (1989).

[16] Tatsumi H, Satoh S. Workstation based 3-D reconstruction system. J Electron Microsc, 41:201 (1992).

[17] Tatsumi H, Satoh S, Takaoki E. Workstation based 3-D reconstruction system: merits and problems in application of ”meta-ball” 3-D reconstruction method to morphological study. Proceedings of the 5th Sapporo International Computer Graphics Symposium, 35-40 (1991).

[18] Fukaya K, Igarashi M, Tatsumi H. The development of color image processing system and its application of region splitting to the stomach cell color image. J Development Policy Studies, 49:155-169 (1992) (in Japanese)

[19] Tatsumi H, Satoh S, Fukaya K, Mochizuki T, Fukuda M. Color image processing for the medical and biological images with NeXT computer. Proceedings of the 6th Sapporo International Computer Graphics Symposium, 85-90 (1992).

[20] Takahashi K, Tatsumi H, Satoh S. Development of education system for hisotology using hyper-multimedea information network: a trial for creating integrated histological image database. Researches Report For Medical Science, 1 ： 7-13(1992) (in Japanese).

[21] Kojima M, Miyake H, Fukuda M, Kumamoto E, Itaya K, Takekawa T, Yamanaka N, Tatsumi H, Kimura K, Yamada H, Katagiri Y. Study on education systems of information science using TCP/IP network system and experiments of its communication at the faculty of medicine Researches Report For Medical Science, 1 ： 1-6(1992)(in Japanese).

[22] Tatsumi H, Nakamura M. Development of a computer-assisted tool for morphological studies: color image handling and processing system. Program of the 4th Japan-Korea Anatomical Joint Meeting Yamagata Seminar :153-154 (1994).

[23] Ishino M, Sasaki T, Fujii N, Miyake H, Tatsumi H, Kojima M, Kimura K, Sakashita M. Establishment of on-line utilization of Bio-Imaging Analyzer and its application to biochemical experiments with protein and nucleic acid. Researches Report For Medical Science, in press (in Japanese).

[24] Tatsumi H. Computer utilization in anatomy: a renaissance in new anatomy. J Clin Exp Med (IGAKU NO AYUMI), 170:in press (1994) (in Japanese).

[25] Tatsumi H. System development for data deposit and retrieval of the JAA using the Internet. Acta Anat Nippon, 68:564-579 (1993) (in Japanese).

[26] Conrad D. A NIC for the Asia Pacific region (APNIC). NFS NETWORK NES, 1:3-11 (1994).